I need some help with the substitution method.
9x +8y= -21
-6x +y=33
I can't figure out the ordered pair.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

I need some help with the substitution method.
9x +8y= -21
-6x +y=33
I can't figure out the ordered pair.
----
Solve the lower equation for "y":
y = 6x+33
---
Substitute into the upper equation:
9x + 8(6x+33) = -21
9x + 48x + 264 = -21
57x = 285
---
x = 5
---
Solve for "y":
y = 6x+33
y = 6*5+33 = 63
===================
Cheers,
Stan H.
============

Answer by solver91311(24713)







Note that the second equation already has one of the variables with a coefficient of 1, namely the . So the simplest strategy would be to solve the second equation for in terms of everything else by adding to both sides:



Now that you have an expression in that is equal to you can substitute back into the first equation:



Now you have a single equation in a single variable that can be solved by ordinary means. Solve for . Then substitute the calculated value of back into the second given equation and solve for the value of . Your ordered pair will be your two calculated values in order .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism